# Spiral of Silence: How is Large Language Model Killing Information Retrieval?—A Case Study on Open Domain Question Answering

### v2

@@ 5.4 Effects of Spiral of Silence on ODQA
(a) Average Query State Transition Number from incorrect to correct (’Average 0->1’) and correct to incorrect (’Average 1->0’) between consecutive iterations, aggregated across all IR methods and datasets.
@@
(b) Ranking of the first retrieved document containing the correct answer in each iteration, with LLM responses being incorrect (EM=0) or correct (EM=1). "First Right From All Sources" refers to the average rank for the first text containing the correct answer, where the source could be either LLM or a human; "First Right From Human" is for human-only sources, both considered across datasets and LLMs.
@@
Figure 7: Effects of "Spiral of Silence" on query and document level of RAG systems.
@@
We will delve into a more nuanced discussion of the impact of the "Spiral of Silence" within the context of ODQA. It is important to note that the influence of the phenomenon is not confined to this scenario; it may also be pertinent across all settings that involve knowledge retrieval, generation, and the influx of text from LLMs. Specifically, our analysis is structured around two dimensions: the query level and the document level.
@@
At the query level, Figure 7(a) signifies the average count of queries shifting between consecutive iterations from incorrect to correct and vice versa, respectively. Notably, during the 1->2 iteration, there is an initial surge in both metrics, which subsequently experience a sharp decline as the iteration count escalates. This suggests that the LLM-generated text’s initial introduction catalyzes a more dynamic state, likely due to the correction of existing errors or the introduction of new inaccuracies. Over time, however, the "Spiral of Silence" effect seems to guide the system towards a state of equilibrium where the transition rate stabilizes to less than 1% per 200 queries. This means most queries maintain their status as either correct or incorrect, indicating that individual query QA results become fixed.
@@
At the document level, we compute the average rank shifts of the first documents containing the correct answer within retrieval results, under different LLM answer states. In Figure 7(b), we observe that: 1) Different Trend of Correct Answer Rankings Based on Source. In instances where EM=0, correct documents from all sources and from humans both tend to be ranked lower over time. When EM=1, the rankings for correct documents from all sources improve slightly, while rankings for correct answers from humans continue to decline. This suggests that the LLM’s correct texts gain prominence in retrieval rankings over time, overshadowing correct texts from human-generated texts. 2) Gradual Dysfunction of the IR System in Incorrect LLM Responses. When the LLM provides incorrect answers (EM=0), there’s a risk that documents that once rose to the top with accurate information might increasingly be obscured by the growing mass of LLM-generated content. This can lead to a scenario where the IR system, originally intended to help users find precise information, becomes less reliable. If it prioritizes and disseminates the LLM’s inaccuracies, a feedback loop could ensue, solidifying these errors. This concerning trend highlights the critical need for ongoing adjustments and improvements to the IR systems to uphold their purpose.
@@ 6 Analysis
To further comprehend the "Spiral of Silence" effect, we illustrate its interaction with misinformation introduced by adversaries using LLMs. Moreover, we test two information filtering mechanisms to alleviate the progression of the effect. For more information on the experimental setup and results, refer to Appendix A.6 and Appendix A.7.
@@ 7 Conclusion
In this study, we initiate our research from empirical observations, aiming to investigate the implications of progressively integrating LLM-generated text into RAG systems. To this end, We employ the ODQA task as a case study to examine both the immediate and extended impacts of LLM text on these systems. Our simulation has revealed the emergence of a "Spiral of Silence" effect, suggesting that without appropriate intervention, human-generated content may progressively diminish its influence within RAG systems. Further investigation into this phenomenon reveals that unchecked accumulation of erroneous LLM-generated information could lead to the overlooking of correct information by IR systems, resulting in harm. We urge the academic community to be vigilant and take measures to prevent the potential misuse of LLM-generated data.
@@
This study aims to present a new perspective on the impact of LLM-generated texts entering the internet on RAG systems. However, the complexity of reality means that it is impossible to account for all variables. The methods of LLM text generation and the mechanisms by which this content enters the retrieval set are constantly changing, which could affect the performance of RAG systems. Dynamic updates to LLMs could also impact the outcomes, and future research should incorporate this variable. While ODQA serves as an insightful approach to evaluate the progression of RAG systems, it is necessary to recognize that ODQA assessments are not exhaustive in capturing the full spectrum of information retrieval scenarios. Nonetheless, the simulation framework proposed in this research is readily adaptable to other tasks that employ RAG systems. Our discussion introduces the "Spiral of Silence" as a potential outcome of the proliferation of LLM-generated texts. Although such a development is not predetermined, given the myriad of factors at play in the real world, this work aims to foster a deeper investigation into the phenomenon and its prospective implications for information diversity in AI-mediated environments.
@@ Appendix A Appendix
A.1 Discussion of Application on "Spiral of Silence"
In aligning the "Spiral of Silence" theory with the focus of this study, emphasis on the aspect of the "individual’s will to express" inherent in the original theory is purposefully diminished. The factors influencing the "Spiral of Silence" phenomenon, as mentioned in  Scheufle and Moy (2000), with media and temporality being the principal elements within RAG systems, directly affect the relative standing of LLM and human texts as the system evolves. While the individual’s desire to express may be indirectly affected by media and temporality, these are not the primary drivers of the "Spiral of Silence" within RAG systems. In RAG systems, we hypothesize that texts generated by LLMs will increasingly be favored in the hierarchy of information retrieval, whereas texts authored by humans might be systematically marginalized, resulting in a structural form of "passive silencing".
@@ A.3 Post-Process Details
  • "I’d be happy to assist you with your question."
  • "According to my knowledge..."
  • "As an AI language model..."
@@
To construct and execute the simulated iterative framework, we adopt a diverse array of tools and technologies to facilitate real-time interaction between various retrieval methods, indexing architectures, and LLMs. We implement the APIs of various LLMs relying on api-for-open-llm [^3] https://github.com/xusenlinzy/api-for-open-llm. With an integration of LangChain [^4] https://github.com/langchain-ai/langchain with Faiss (Johnson et al., 2019) and Elasticsearch [^5] https://github.com/elastic/elasticsearch, we execute batched incremental updates of LLM-generated documents in each iteration, thus simulating the process of document index updating by search engines in real-world scenarios. To maintain the diversity of the generated texts, we set the temperature at 0.7 for all LLMs. In each iteration of the experiment, except for the zero-shot setting, we keep the size of the context document set $D_{i}^{\prime}$ fixed at 5. We rerank the first 100 documents recalled by the retrieval method when the step is applied. We apply the LLMs to generate response text, post-process via rules, and then merge their outputs into the index for each query in every iteration. Therefore, for each iteration, we will add 4k new samples to the index. The total number of iterations $t$ is set at 10, which results in a total of 40k invocations of the LLMs for each experimental run. We will make our code and data publicly available for further research.
@@ A.5 Results on WebQ and TriviaQA
Figure 9: Average Percentage of texts from various sources within the top 50 search results over multiple iterations across different search and methods.
@@
Figure 10: 3-gram Self-BLEU score for the top 5 search results over iterations, from the original dataset (Ori.) to the inclusion and subsequent iterations with LLM-generated texts.
A.6 Misinformation in the Iteration
@@
In the previous sections, we explored the impact of non-maliciously generated texts by LLMs on the evolution of the RAG system over time. In this section, we will discuss the persistence of the "Spiral of Silence" effect when attackers deliberately inject specific misinformation into the RAG system, how misinformation could affect the RAG system over time, and the feasibility of targeted misinformation injection.
@@
Table 5: Percentage of LLM-generated documents with Misinformation occupying the top-5 retrieval results, after augmenting each query with five documents generated by LLMs. Data entries framed by a blue background indicate a majority presence of human-generated documents, while entries with a purple background denote a predominance of LLM-generated documents.
Analysis of Experimental Results: From the experimental results, we can observe that: 1) The "Spiral of Silence" still exists. We first investigate the presence of the "Spiral of Silence" phenomenon when misinformation targeting the objective is injected into the corpus. As shown in Table 5, although the majority of the injected information is misleading, the content generated by the LLMs is still quickly ranked at the top by the retrieval systems, taking a dominant position. When comparing four different retrieval methods, the BM25 algorithm shows greater robustness than the others, being least affected by the LLM-generated content. However, it is noteworthy that approximately 20% of the content generated by the LLM could still be quickly placed in the forefront of the search results by the BM25 algorithm.
@@
Figure 12: Correlation between the number of top 5 search results containing the correct answer ("Context Right Num") and the accuracy of responses given by LLMs on the NQ dataset when adding Misinformation. The responses are categorized based on EMllm score: EM_L=1 for correct and EM_L=0 for incorrect. The overall number of queries that the LLMs answered correctly (EM_L=1 Total) and incorrectly (EM_L=0 Total), along with the average retrieval accuracy (Acc@5) are shown by dashed lines. The results are averaged across different LLMs, retrieval and ranking methods.
Figure 11 illustrates that over time, human-written texts are gradually excluded from the searchable range, and as depicted in Figure 12, the phenomenon of homogenization of opinions in search results persists. This further indicates that regardless of the accuracy of the LLM-generated information, the spiral of silence phenomenon remains present. 2) The RAG system has a limited degree of self-correction capability. LLM-generated texts containing misinformation lead to a significant decline in retrieval and QA performance based on the RAG system compared to results in Section 5.2. With the continuation of the iteration process, the number of correct answers increased and the number of incorrect answers decreased, albeit by a small margin. This suggests that the RAG system has a certain degree of self-correction capability, which may stem from the model’s own knowledge or the human-written texts containing correct information retrieved in the initial stages.
@@
3) The introduction of a small amount of the LLM-generated texts with specific misleading information during the iterative process could inject such information into the RAG output. In Figure 13, we quantify the $\text{EM}_{llm}$ metric of the original and specific misleading answers generated by the RAG system based on four retrieval methods at various iteration stages. The results show that after the purposeful addition of misleading information (before the first iteration), the proportion of RAG system-generated answers containing specific misleading information significantly increases, especially on the NQ and PopQA datasets, where the proportion of incorrect answers exceeds that of correct ones, and the influence of misleading answers persisted over time. However, the BM25 algorithm exhibits relatively higher robustness, and the $\text{EM}_{llm}$ of incorrect answers output by the RAG system based on it remains lower than the other three retrieval methods. The experimental results of this section reveal that despite the presence of self-correcting mechanisms, the injection of specific misleading information can still severely compromise the system’s accuracy and enable the manipulation of the RAG system to consistently output specific misinformation in response to certain questions. Therefore, without timely intervention, the spiral of silence phenomenon could marginalize accurate information, leading to severe misinformation consequences.
A.7 Attempts to Alleviate the Spiral of Silence
The "Spiral of Silence" effect could lead to the marginalization of human-generated text expression and further enhance the homogeneity of retrieval outcomes. If left unaddressed, this phenomenon could precipitate a series of adverse repercussions. To mitigate or eliminate the influence of the "Spiral of Silence" effect, this section initiates a discussion on two fronts. First, from the perspective of the authenticity of sources, we employ the widely used AIGC detection technologies to filter out and exclude all non-human-produced texts at the top of the search results. Second, addressing the validity of content, we strive to maintain diversity among the top search results to overcome potential issues caused by excessive homogenization.

### v3

@@ 5.4 Effects of “Spiral of Silence” on ODQA
(a) Average Query State Transition Number from incorrect to correct (“Average 0->1”) and correct to incorrect (“Average 1->0”) between consecutive iterations, aggregated across all IR methods and datasets.
@@
(b) Ranking of the first retrieved document containing the correct answer in each iteration, with LLM responses being incorrect (EM=0) or correct (EM=1). “First Right From All Sources” refers to the average rank for the first text containing the correct answer, where the source could be either LLM or human; “First Right From Human” is for human-only sources, both considered across datasets and LLMs.
@@
Figure 7: Effects of “Spiral of Silence” on query and document level of RAG systems.
@@
We will delve into a more nuanced discussion of the impact of the “Spiral of Silence” within the context of ODQA. It is important to note that the influence of the phenomenon is not confined to this scenario; it may also be pertinent across all settings that involve knowledge retrieval, generation, and the influx of text from LLMs. Specifically, our analysis is structured around two dimensions: the query level and the document level.
@@
At the query level, Figure 7(a) signifies the average count of queries shifting between consecutive iterations from incorrect to correct and vice versa, respectively. Notably, during the 1->2 iteration, there is an initial surge in both metrics, which subsequently experience a sharp decline as the iterations continue. This suggests that the LLM-generated text’s initial introduction catalyzes a more dynamic state, likely due to the correction of existing errors or the introduction of new inaccuracies. Over time, however, the “Spiral of Silence” effect seems to guide the system towards a state of equilibrium where the transition rate stabilizes to less than 1% per 200 queries. This means most queries maintain their status as either correct or incorrect, indicating that individual query QA results become fixed.
@@
At the document level, we compute the average rank shifts of the first documents containing the correct answer within retrieval results, under different LLM answer states. In Figure 7(b), we observe that: 1) Different Trends of Correct Answer Rankings Based on Source: In instances where EM=0, correct documents from all sources (“First Right From ALL Sources”) and from humans (“First Right From Human") both tend to be ranked lower over time. When EM=1, the rankings for correct documents from all sources improve slightly, while rankings for correct answers from humans continue to decline. This suggests that the LLM’s correct texts gain prominence in retrieval rankings over time, overshadowing correct texts from human-generated texts. 2) Gradual Dysfunction of the IR System in Incorrect LLM Responses: When the LLM provides incorrect answers (EM=0), there’s a risk that documents that once rose to the top with accurate information might increasingly be obscured by the growing mass of LLM-generated content. This can lead to a scenario where the IR system, originally intended to help users find precise information, becomes less reliable. If it prioritizes and disseminates the LLM’s inaccuracies, a feedback loop could ensue, solidifying these errors. This concerning trend highlights the critical need for ongoing adjustments and improvements to the IR systems to uphold their purpose.
@@ 6 Analysis
To further comprehend the “Spiral of Silence” effect, we illustrate its interaction with misinformation introduced by adversaries using LLMs. Moreover, we test two information filtering mechanisms to alleviate the progression of the effect. For more information on the experimental setup and results, refer to Appendix A.7 and Appendix A.8.
@@ 7 Conclusion
In this study, we initiate our research from empirical observations, aiming to investigate the implications of progressively integrating LLM-generated text into RAG systems. To this end, We employ the ODQA task as a case study to examine both the immediate and extended impacts of LLM text on these systems. Our simulation has revealed the emergence of a “Spiral of Silence” effect, suggesting that without appropriate intervention, human-generated content may progressively diminish its influence within RAG systems. Further investigation into this phenomenon reveals that unchecked accumulation of erroneous LLM-generated information could lead to the overlooking of correct information by IR systems, resulting in harm. We urge the academic community to be vigilant and take measures to prevent the potential misuse of LLM-generated data.
@@
This study aims to present a new perspective on the impact of LLM-generated texts entering the internet on RAG systems. However, the complexity of reality means that it is impossible to account for all variables. The methods of LLM text generation and the mechanisms by which this content enters the retrieval set are constantly changing, which could affect the performance of RAG systems. While ODQA serves as an insightful approach to evaluate the progression of RAG systems, it is necessary to recognize that ODQA assessments are not exhaustive in capturing the full spectrum of information retrieval scenarios. Nonetheless, the simulation framework proposed in this research is readily adaptable to other tasks that employ RAG systems. Our discussion introduces the “Spiral of Silence” as a potential outcome of the proliferation of LLM-generated texts. Although such a development is not predetermined, given the myriad of factors at play in the real world, this work aims to foster a deeper investigation into the phenomenon and its prospective implications for information diversity in AI-mediated environments.
@@
Acknowledgements
This work is supported by the National Natural Science Foundation of China (62122077/62272439/62106251), Beijing Municipal Science and Technology Project (No. Z231100010323002), and the Fundamental Research Funds for the Central Universities.
@@ Appendix A Appendix
A.1 Discussion of Application on “Spiral of Silence”
In aligning the “Spiral of Silence” theory with the focus of this study, emphasis on the aspect of the “individual’s will to express” inherent in the original theory is purposefully diminished. The factors influencing the “Spiral of Silence” phenomenon, as mentioned in  Scheufle and Moy (2000), with media and temporality being the principal elements within RAG systems, directly affect the relative standing of LLM and human texts as the system evolves. While the individual’s desire to express may be indirectly affected by media and temporality, these are not the primary drivers of the “Spiral of Silence” within RAG systems. In RAG systems, we hypothesize that texts generated by LLMs will increasingly be favored in the hierarchy of information retrieval, whereas texts authored by humans might be systematically marginalized, resulting in a structural form of “passive silencing".
@@ A.3 Post-Process Details
  • ‘‘I’d be happy to assist you with your question.’’
  • ‘‘According to my knowledge...’’
  • ‘‘As an AI language model...’’
@@
To construct and execute the simulated iterative framework, we adopt a diverse array of tools and technologies to facilitate real-time interaction between various retrieval methods, indexing architectures, and LLMs. We implement the APIs of various LLMs relying on api-for-open-llm [^3] https://github.com/xusenlinzy/api-for-open-llm. With integration of LangChain [^4] https://github.com/langchain-ai/langchain with Faiss (Johnson et al., 2019) and Elasticsearch [^5] https://github.com/elastic/elasticsearch, we execute batched incremental updates of LLM-generated documents in each iteration, thus simulating the process of document index updating by search engines in real-world scenarios. To maintain the diversity of the generated texts, we set the temperature at 0.7 for all LLMs. In each iteration of the experiment, except for the zero-shot setting, we keep the size of the context document set $D_{i}^{\prime}$ fixed at 5. We rerank the first 100 documents recalled by the retrieval method when the step is applied. We apply the LLMs to generate response text, post-process via rules, and then merge their outputs into the index for each query in every iteration. Therefore, for each iteration, we will add 4k new samples to the index. The total number of iterations $t$ is set at 10, which results in a total of 40k invocations of the LLMs for each experimental run.
@@ A.5 Results on WebQ and TriviaQA
Figure 8: Long-Term RAG performance for WebQ and TriviaQA. The upper section illustrates the retrieval outcomes for various methods, while the lower section depicts the average EM across LLMs. Iteration 1 represents the results following the incorporation of zero-shot LLM-generated text. Abbreviated re-ranking methods in the legend are: +U for UPR, +M for MonoT5, and +BR for BGE-Reranker.
@@
Figure 9: Average Percentage of texts from various sources within the top 50 search results over multiple iterations across different search and methods for WebQ and TriviaQA.
@@
Figure 10: 3-gram Self-BLEU score for the top 5 search results over iterations for WebQ and TriviaQA, from the original dataset (Ori.) to subsequent iterations including LLM-generated texts.
A.6 Simulation with Model Evolution
In this section, we explore the impact of the continuous influx of text generated by evolving LLMs on RAG systems over time. To this end, we experiment by using a series of LLMs, ranging from weak to strong, to observe the phenomenon.
Experimental Setup: We conduct experiments based on BM25 on the NQ and WebQ datasets. For the text generation models, we use progressively larger versions of the Qwen1.5 model (Bai et al., 2023) across different iterations: 0.5b, 1.8b, and 4b versions for the first three iterations. For iterations four to seven, we employ Qwen1.5-7b-Chat (Bai et al., 2023), LLaMA2-7b-Chat (Touvron et al., 2023), and Baichuan2-7b-Chat (Yang et al., 2023). Finally, for the eighth to tenth iterations, we utilize GPT-3.5-Turbo (OpenAI, 2022), LLaMA2-13b-Chat (Touvron et al., 2023), and Qwen1.5-14b-Chat (Bai et al., 2023). This setup simulates the impact of increasing model capabilities on the RAG system over time.
Dataset NQ WebQ
Iteration Acc@20 EM Acc@20 EM
1 68.5 29.5 66.0 35.5
2 67.5 28.5 65.0 38.7
3 64.0 29.3 63.5 38.0
4 63.0 32.5 62.5 43.5
5 60.5 32.3 62.5 41.8
6 58.0 32.0 61.5 43.3
7 55.5 30.7 61.5 43.2
8 52.0 33.7 60.5 44.0
9 51.5 32.8 61.0 44.3
10 47.5 33.7 58.5 44.2
Table 3: BM25 retrieval performance (Acc@20) and QA performance (EM) across iterations for NQ and WebQ under the evolving LLM setting.
Iteration NQ WebQ
1 77.7 87.0
2 45.9 64.3
3 27.2 48.6
4 18.0 40.2
5 14.6 36.2
6 12.3 32.6
7 10.1 28.9
8 8.6 25.6
9 6.7 21.9
10 5.8 19.2
Table 4: Percentage of human text in the top 5 BM25 retrieval results across iterations for NQ and WebQ under the evolving LLM setting.
Model Name NQ WebQ
Qwen1.5-0.5b-Chat 12.5 6.9
Qwen1.5-1.8b-Chat 7.4 4.4
Qwen1.5-4b-Chat 8.5 6.3
Qwen-7b-Chat 1.1 1.7
LLaMA2-7b-Chat 20.9 18.5
Baichuan2-7b-Chat 10.3 7.6
LLaMA2-13B-Chat 19.1 19.9
Qwen1.5-14b-chat 0.9 1.8
GPT-3.5-Turbo 13.5 13.7
Human 5.8 19.2
Table 5: Percentage of generated text by different LLMs in the top 5 BM25 retrieval results at the end of the simulation (Iteration 10) for NQ and WebQ under the evolving LLM setting.
Figure 11: Correlation between the number of top 5 BM25 search results containing the correct answer ("Context Right Num") and the accuracy of responses given by evolving LLMs on the NQ dataset. The responses are categorized based on n Exact Match (EM) score: EM=1 for correct and EM=0 for incorrect. The overall number of queries that the LLMs answered correctly (EM=1 Total) and incorrectly (EM=0 Total), along with the average retrieval accuracy (Acc@5) are shown by dashed lines. The results are averaged across different LLMs.
Iteration NQ WebQ
1 27.5 25.4
2 49.8 39.5
3 65.6 53.3
4 74.4 60.9
5 70.8 59.2
6 72.3 60.9
7 74.8 63.6
8 75.5 64.5
9 75.3 64.7
10 78.2 68.4
Table 6: 3-gram Self-BLEU score for the top 5 BM25 search results over iterations for NQ and WebQ under the evolving LLM setting.
Analysis of Experimental Results: From the experimental results, we can observe that: 1) Retrieval Performance Continues to Decline: According to the Acc@20 results reported in Table 3, even with the gradual improvement of LLM performance in the simulation, the retrieval performance exhibits a general downward trend over time. Compared to the results in Table 1, the improvements in retrieval performance are also limited for weaker LLMs in the initial stage (73.5 $\rightarrow$ 68.5, 71.0 $\rightarrow$ 66.0). 2) QA Results Show Improvement: Based on the EM results reported in Table 3, we can see that with the expansion of the LLM scale, the QA results have improved, especially from the third to the fourth iteration, where the model expanded from less than 4B to 7B, resulting in more than a 10% improvement in QA results on both datasets. Considering that QA results remain stable in Figure 3 when LLMs are unchanged, the current improvement likely stems from the enhanced capabilities of the LLM itself. 3) The Existence of the “Spiral of Silence” Phenomenon: In Table 4, we present the proportion of human text in the top 5 retrieval results during the iterations. It can be observed that LLM-generated text still dominates the retrieval results, and the influence of human text continues to decline. In Table 5, we report the proportion of different LLM-generated texts in the top 5 retrieval results after the tenth iteration. Overall, texts generated by larger-scale LLMs occupy a slightly higher proportion, but this is not absolute. Table 6 presents the diversity of retrieval results, while Figure 11 illustrates the relationship between retrieval accuracy and the generated answers. As the model evolves, the experimental results remain consistent with the conclusion in Section 5.3: the diversity of retrieval results continues to decline, and polarization occurs. In summary, the experimental results indicate that the evolution of LLMs will not prevent the occurrence of the “Spiral of Silence” phenomenon.
A.7 Effects of Misinformation
@@
In previous sections, we explored the impact of non-maliciously LLM-generated texts on the evolution of the RAG system over time. In this section, we will discuss the persistence of the “Spiral of Silence” when attackers deliberately inject specific misinformation into the RAG system, how misinformation could affect the system over time, and the feasibility of targeted misinformation injection.
@@
Table 9: Percentage of LLM-generated documents with Misinformation occupying the top 5 retrieval results, after augmenting each query with five documents generated by LLMs. Data entries framed by a blue background indicate a majority presence of human-generated documents, while entries with a purple background denote a predominance of LLM-generated documents.
Analysis of Experimental Results: From the experimental results, we can observe that: 1) The “Spiral of Silence” Still Exists: We first investigate the presence of the “Spiral of Silence” phenomenon when misinformation targeting the objective is injected into the corpus. As shown in Table 9, although the majority of the injected information is misleading, the content generated by the LLMs is still quickly ranked at the top by the retrieval systems, taking a dominant position. When comparing four different retrieval methods, the BM25 algorithm shows greater robustness than the others, being least affected by the LLM-generated content. However, it is noteworthy that approximately 20% of the content generated by the LLM could still be quickly placed in the forefront of the search results by the BM25 algorithm.
@@
Figure 13: Correlation between the number of top 5 search results containing the correct answer (“Context Right Num”) and the accuracy of responses given by LLMs on the NQ dataset when adding Misinformation. The responses are categorized based on EMllm score: EM_L=1 for correct and EM_L=0 for incorrect. The overall number of queries that the LLMs answered correctly (EM_L=1 Total) and incorrectly (EM_L=0 Total), along with the average retrieval accuracy (Acc@5) are shown by dashed lines. The results are averaged across different LLMs, retrieval and ranking methods.
Figure 12 illustrates that over time, human-written texts are gradually excluded from the searchable range, and as depicted in Figure 13, the phenomenon of homogenization of opinions in search results persists. This further indicates that regardless of the accuracy of the LLM-generated information, the “Spiral of Silence” phenomenon remains present. 2) The RAG System has a Limited Degree of Self-Correction Capability: LLM-generated texts containing misinformation lead to a significant decline in retrieval and QA performance based on the RAG system compared to results in Section 5.2. With the continuation of the iteration process, the number of correct answers increased and the number of incorrect answers decreased, albeit by a small margin. This suggests that the RAG system has a certain degree of self-correction capability, which may stem from the model’s knowledge or the human-written texts containing correct information retrieved in the initial stages.
@@
3) The Introduction of a Small Amount of the LLM-generated Texts with Specific Misleading Information during the Iterative Process could Inject such Information into the RAG Output: In Figure 14, we quantify the $\text{EM}_{llm}$ metric of the original and specific misleading answers generated by the RAG system based on four retrieval methods at various iteration stages. The results show that after the purposeful addition of misleading information (before the first iteration), the proportion of RAG system-generated answers containing specific misleading information significantly increases, especially on NQ and PopQA, where the proportion of incorrect answers exceeds that of correct ones, and the influence of misleading answers persisted over time. However, the BM25 algorithm exhibits relatively higher robustness, and the $\text{EM}_{llm}$ of incorrect answers output by the RAG system based on it remains lower than the other three retrieval methods. The experimental results of this section reveal that despite the presence of self-correcting mechanisms, the injection of specific misleading information can still severely compromise the system’s accuracy and enable the manipulation of the RAG system to consistently output specific misinformation in response to certain questions. Therefore, without timely intervention, the “Spiral of Silence” phenomenon could marginalize accurate information, leading to severe misinformation consequences.
A.8 Attempts to Alleviate the “Spiral of Silence"
The “Spiral of Silence” effect could lead to the marginalization of human-generated text expression and further enhance the homogeneity of retrieval outcomes. If left unaddressed, this phenomenon could precipitate a series of adverse repercussions. To mitigate or eliminate the influence of the “Spiral of Silence” effect, this section initiates a discussion on two fronts. First, from the perspective of the authenticity of sources, we employ the widely used AIGC detection technologies to filter out and exclude all non-human-produced texts at the top of the search results. Second, addressing the validity of content, we strive to maintain diversity among the top search results to overcome potential issues caused by excessive homogenization.